OpenStudy (anonymous):

#1: How many meters are in 0.05 kilometers? A. 50 B. 500 C. 5,000 D. 50,000 **MY answer: A is that right?? :)

OpenStudy (anonymous):

#2: A container has the dimensions of 30 cm x 50 mm x 0.1 m. The density of its contents is 2.5 g/cm^3. What is the mass of the substance in kilograms? (D = m/V) A. 3.8 kg B. 38 kg C. 380 kg D. 3,800 kg ***Idk :( please explain? :) thanks :)

OpenStudy (anonymous):

heyy!! :) so did i get #1 right?? is it A. 50??

OpenStudy (aaronq):

hey ! yep # 1 is A for #2 use the formula Density = mass/ Volume

OpenStudy (aaronq):

1 cm^3 = 1 mL

OpenStudy (aaronq):

do find the volume of the box.. length x width x height

OpenStudy (anonymous):

wait what?? so i use mass/volume=d so how do i find the mass?? and the density is 2.5 g/cm^3? so i plug in and get this? 2.5 g/cm^3=m/?? idk how to find the volume :( is this the lxwxh?? 30 cm x 50 mm x 0.1 m.

OpenStudy (aaronq):

haha yep, to get the volume of a rectangular box you multiply all 3 lengths in the 3 dimensions

OpenStudy (aaronq):

but you have to convert all the units to the same unit

OpenStudy (anonymous):

how do i do that?? I'm so bad at converting :(

OpenStudy (aaronq):

1 m = 100 cm = 1000 mm

OpenStudy (aaronq):

it's easy with the metric system, i hate inches and yards all that nonsense

OpenStudy (anonymous):

so i get this? 30 cm x 50 mm x 0.1 m =0.3m x 0.5m x 0.1 m ??

OpenStudy (aaronq):

30 cm = 0.3 m 50 mm= 0.05 m remember just move the decimal point over 3 times when dividing by 1000

OpenStudy (anonymous):

ohh so i multiply all that and get this? 0.0015 m ???

OpenStudy (aaronq):

yep m^3 though remember your units now notice that they give you density in units of g/cm^3 so you have to convert either one to match .. id suggest convert m^3 to cm^3 ..

OpenStudy (anonymous):

wait so i get this? V=30cm x 5cm x 10cm=1500 cm3

OpenStudy (anonymous):

???

OpenStudy (anonymous):

and then i get this?? d=m/V which is the same as m=dV so i get this??? m=2.5g/cm3 x 1500cm3 = 3750g = 3.75kg which is about 3.8. So the answer is 3.8kg??? is answer A. 3.8 kg my answer?? @aaronq ?? :D

OpenStudy (aaronq):

yep!

OpenStudy (aaronq):

sorry my internet connection is garbage and i keep cutting off

OpenStudy (aaronq):

but you are correct ma friend

OpenStudy (anonymous):

oh it's okay!! YAYYYY :) woot woot!! Thanks :)

OpenStudy (anonymous):

and yeah my internet is like sooo annoying sometimes :(

OpenStudy (aaronq):

wooo (y) no problem, it was all you man!

OpenStudy (aaronq):

yeah! any other site would load except this one, it's silly

OpenStudy (anonymous):

haha thanks :) and yeah OS is supeeeerr slow sometimes, like it crashes a lot and lags too :(

OpenStudy (anonymous):

haha thanks :) and yeah OS is supeeeerr slow sometimes, like it crashes a lot and lags too :(